The primitive translation vectors of a crystal lattice are given by :
a₁=(√(3)a/2)i + (a/2)j
a₂=-(√(3)a/2)i + (a/2)j
a₃= ck
To calculate the volume of the primitive cell and obtain the primitive translation vectors of the reciprocal lattice.
1
Expert's answer
2019-11-27T10:16:23-0500

Volume of primitive cell is given by

"V=a1.(a2\times a3)"


"a2\times a3=\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 a & b & c \\ \hline
 d & e & f \\
 \hdashline
 g & h & i
\end{array}"

Where a=i, b=j, c=k

From vector a2 and a3, we have

d= -(√(3)a/2), e=(a/2), f=0

g=0, h=0, i=c

Now solve for determinants, we get cross product of a2 and a3=a(ei-fh)-b(di-fg)+c(dh-eg)

=a(ei)-b(di)

=i(a/2xc)+j(√(3)a/2xc)

Volume of primitive cell

V=a1.(a2xa3)={(√(3)a/2)i + (a/2)j}.{(a/2xc)i+(√(3)a/2xc)j}

V=a2c√3/2

And

the primitive translation vectors of the reciprocal lattice is given by

"b1=(2\pi/a√3)i+(2\pi/a)j"

"b2=(-2\pi/a√3)i+(2\pi/a)j"

"b3=(2\pi/c)k"